OpenStudy (anonymous):

32+1=2-2 help me find the solution please

OpenStudy (king):

dude first if the above thing is true then maths will become sriously easy bcuz all nos. will be equal to 0! bcuz 2-2=0 33=0 !

OpenStudy (anonymous):

so the solution is 0

OpenStudy (compassionate):

*Facepalm.* No Cdel.

OpenStudy (anonymous):

what do you mean

OpenStudy (compassionate):

Does 33 = 0?

OpenStudy (anonymous):

33 what

OpenStudy (compassionate):

ohmyisheserious.jpg 33 is a number. 33 ones.

OpenStudy (anonymous):

ok this is my first time taking math I do not need negative comments

OpenStudy (compassionate):

Okay, are you aware of what the number 33 is?

OpenStudy (anonymous):

the solution

OpenStudy (anonymous):

32+1=33

OpenStudy (compassionate):

33 = 33, 33 cannot equal zero.

OpenStudy (anonymous):

so is the solution 33

OpenStudy (king):

dude 33 is a NUMBER 0 is also a NUMBER one NUMBER CANNOT EQUAL ANOTHER NUMBER UNTIL AND UNLESS THEY ARE THE SAME..dont ask why bcuz thatz hw maths is ..its like telling all NUMBERS ARE EQUAL TO 0!

OpenStudy (anonymous):

ok

OpenStudy (compassionate):

We've covered some ground here.

OpenStudy (anonymous):

I still do not know the solution

OpenStudy (king):

DUDE EITHER U ARE MESSIN AROUND WITH US OR URE COMPLETELY INSANE/DUMB!

OpenStudy (anonymous):

I am not insane I am asking for help and you are a idot for all the negative comments you have said you do not need to be a tutor.

OpenStudy (asnaseer):

I think we all need to calm down here. I am sure cdelomas is not messing about but just trying to get help. @cdelomas - have you stated your question correctly? is it: 32 + 1 = 2 - 2 or: 32 + L = 2 - 2 or something else?

OpenStudy (anonymous):

it is the first one

OpenStudy (king):

see..!@asnaseer

OpenStudy (asnaseer):

if it is the first one, then you are stating that "32 + 1" is the same thing as "2 - 2" i.e. you are stating that "33" is the same thing as "0" are you sure that is what the question states?

OpenStudy (asnaseer):

@king please calm down - lets try and help here

OpenStudy (anonymous):

that is the question

OpenStudy (asnaseer):

where did you get this question from?

mathslover (mathslover):

hmn I think @cdelomas needs something practical to know about : @cdelomas are 33 chocolates equal to 2 chocolates?

OpenStudy (anonymous):

Would the solution possibly be that \[33 \neq 2\] ?

OpenStudy (asnaseer):

or is the question asking you state whether this is true or false?

OpenStudy (anonymous):

no

OpenStudy (king):

@cshalvey @asnaseer yeah both are possible...

OpenStudy (asnaseer):

@cdelomas is this a homework question or is it out of a book?

OpenStudy (anonymous):

it is from a study plan question

mathslover (mathslover):